# Walk Wisely on Graph: Knowledge Graph Reasoning with Dual Agents via Efficient Guidance-Exploration

### v1

@@ Introduction
$\text{where}\ \Delta(s_{t}^{c},s_{t+1}^{c})=D(s_{t}^{c},s_{\text{target}}^{c})-D(s_{t+1}^{c},s_{\text{target}}^{c})$
$s_{t+1}^{c}$ comes from the next state generated by the policy network. In contrast to the default reward, dynamic path feedback uses the reward function to score the GIANT’s path in a rollout rather than simply identifying whether it has reached the correct target cluster. Even if GIANT does not reach the correct target cluster in a rollout, it will evaluate the quality of the path, thereby accelerating the learning process. In Appendix D, we prove that GIANT learns optimal policy in dynamic path feedback is consistent with default rewards.
Algorithm 1 FULORA Training Algorithm
0:    KG $\mathcal{G}$; Initial policy networks parameters $\theta_{\pi^{e}}$ and $\theta_{\pi^{c}}$; Initial Lagrange multiplier parameter $\theta_{\lambda}$ ;Source entity and cluster nodes $e_{s}$ and $c_{s}$; Entity-level query $r_{q}$; Target entity and cluster nodes $e_{o}$ and $c_{o}$; Maximum path length $T$
0:    parameters $\theta_{\pi^{e}}$, $\theta_{\pi^{c}}$, $\theta_{\lambda}$
1:  for $t=0,...,T-1$ do
2:     Set default cluster-level reward $r_{c}=1$ if $c_{t}=c_{o}$ otherwise $r_{c}=0$
3:     Set default entity-level reward $r_{e}=1$ if $e_{t}=e_{o}$ otherwise $r_{e}=0$
5:     Compute $J(\theta_{\pi^{c}})$, $J(\theta_{\pi^{e}})$, $J(\theta_{\lambda})$ based on Eq. (10)-(12).
6:  end for
8:  return  $\theta_{\pi^{e}}$, $\theta_{\pi^{c}}$, $\theta_{\lambda}$
@@
Dataset #Ent #Rel #Fact #Que #$\textbf{\#Deg}_{\textbf{mean}}#$$\textbf{\#Deg}_{\textbf{med}}#$
@@
Table 1: The statistics of some benchmark KG datasets. #Deg is the outgoing degree of every entity that can indicate the sparsity level.
As can be seen from the statistical indicators, these three datasets represent standard KG, sparse KG and Dense KG respectively. In our selection of the baseline, we are not only evaluating it against the state-of-the-art multi-hop reasoning methods (DeepPath (Xiong, Hoang, and Wang 2017), MINERVA (Das et al. 2018), M-Walk (Shen et al. 2018), AttnPath (Wang et al. 2019), SQUIRE (Bai et al. 2022) and CURL (Zhang et al. 2022)), but also against with other embedding-based KG reasoning methods (TransE (Bordes et al. 2013), DistMult (Yang et al. 2015), ComplEx (Trouillon et al. 2016) and LMKE (Wang et al. 2022)). FULORA and all baselines are implemented under the Pytorch framework and run on the NVIDIA 3080Ti GPU. We set the default path length to 3, and all the results are the average of the results in five experiments.
Link Prediction Results
For each triplet $(e_{s},r_{q},e_{o})$ in the test set, we convert it to a triplet query $(e_{s},r_{q},?)$, and then use embedding-based models or multi-hop based models and a beam search width of 50 to get the ranking list of tail entity. Following the previous work (Bordes et al. 2013), here we use two metrics: (1) the mean reciprocal rank of all correct tail entities (MRR), and (2) the proportion of correct tail entities ranking in the top K (Hits@K) for evaluation.
@@
TransE 51.4 45.6 67.8 75.1 35.9 28.9 46.4 53.4 36.1 24.8 40.1 45.0
DistMult 68.0 61.0 73.3 79.5 43.3 41.0 44.1 47.5 37.0 27.5 41.7 56.8
ComplEx 68.4 61.2 76.1 82.1 41.5 38.2 43.3 48.0 39.4 30.3 43.4 57.2
LMKE 74.6 71.7 84.7 89.5 53.6 45.1 66.0 79.4 41.2 31.8 46.2 56.9
MINERVA 67.5 58.8 74.6 81.3 44.8 41.3 45.6 51.3 27.1 19.2 30.7 42.6
M-Walk 70.7 63.2 75.7 81.9 43.7 41.5 44.7 54.3 23.4 16.8 24.5 40.3
AttnPath 69.3 62.7 73.9 80.1 42.9 40.7 44.3 52.9 31.9 24.1 40.4 43.8
SQUIRE 71.1 68.2 81.4 87.2 48.2 45.0 51.4 59.7 37.4 28.0 42.1 52.5
CURL 70.8 66.7 78.6 84.3 46.0 42.9 47.1 52.3 30.6 23.9 38.1 50.9
FULORA 72.5 69.4 79.7 89.2 49.1 45.6 50.7 57.2 33.4 25.4 39.2 50.7
@@
Table 2: Link prediction results on three datasets from NELL, WordNet and Freebase. All metrics are multiplied by 100. The best score of multi-hop reasoning models is in bold, and the best score of embedding-based models is underlined.
As shown in Table 2, we first display the performances of FULORA and all baselines on NELL-995, WN18RR and FB15K-237. From the table, we can observe that our model outperforms all previous multi-hop reasoning models on standard KG (NELL-995) and sparse KG (WN18RR). Especially on NELL-995, where our model gains marked improvements compared with the best multi-hop reasoning baseline SQUIRE and dual-agent model CURL. In addition, FULORA scores similarly to the best embedding-based model LMKE. Since FB15K-237 contains much larger number of 1-to-M than the M-to-1 relation instances (Wan et al. 2020; Zhang et al. 2022), multi-hop reasoning methods is prone to be stuck in the local entity nodes with high-degree centrality, renders it hard to reach the correct entity. Compared with strigent multi-hop reasoning models (MINERVA, M-Walk and CURL), FULORA’s attention mechanism alleviates this issue by forcing the agent to focus on the neighbor which is highly relative to the query. After averaging results on three datasets, we observe that FULORA leads to overall improvements relative to the dual-agent model (CURL) by 2.5%, 2.3%, 1.3%, 3.2%, in terms of MRR and Hits@1, 3, 10.
Fact Prediction Results
As opposed to link prediction, fact prediction task is concerned with verifying the veracity of an unknown fact, the true test triplets are ranked with some generated false triplets. Since we share a similar query-answering mechanism as CURL (Zhang et al. 2022), FULORA is capable of identifying the most appropriate entity for a given query and eliminates the need to evaluate negative samples of any particular relation. In the experiments, the dual agents try to infer and walk through the cluster-level and entity-level respectively to reach the correct target under removing all links of groundtruth relations in the original KG. Here, we report Mean Average Precision (MAP) scores for various relation tasks of NELL-995. We reuse the results of TransE, TransR, PRA, DeepPath, MINERVA, M-Walk on seven tasks already reported in (Zhang et al. 2022). As demonstrated in Table 3, FULORA produces a satisfying result in most tasks, contributing an average gain of 9.1% relative to the multi-hop based reasoning approaches (PRA, DeepPath, MINERVA, M-Walk and CURL) and 16.1% gain compared to the embedding-based approaches (TransE and TransR).
@@
Recall the motivation of FULORA, we tend to address the issue of multi-hop reasoning models being unable to infer correct answer due to the lack of short direct path by improving long-distance reasoning ability. This issue is significant on standard KG and sparse KG, so we conduct the following experiments on NELL-995 and WN18RR. For effective evaluation, we compare our model with MINERVA and CURL in NELL-995 and WN18RR, where we remove the most frequently-visited short paths found by the bi-directional search (Xiong, Hoang, and Wang 2017) inside KGs.
Task TransE TransR PRA DeepPath MINERVA M-Walk CURL FULORA
PersonBornInLocation 62.7 67.3 54.7 75.1 80.0 84.7 82.7 84.4
OrgHeadquarteredInCity 62.0 65.7 81.1 79.0 94.0 94.3 94.8 95.1
AthletePlaysForTeam 62.7 67.3 54.7 75.0 80.3 84.7 82.9 86.0
AthletePlaysInLeague 77.3 91.2 84.1 96.0 94.2 96.1 97.1 97.3
AthletePlaysSport 87.6 96.3 47.4 95.7 98.0 98.3 98.4 98.5
TeamPlaysSport 76.1 81.4 79.1 73.8 88.0 88.4 88.7 90.2
WorksFor 67.7 69.2 68.1 71.1 81.0 83.2 82.1 84.3
@@
Cluster State Similarity.   One of the most significant contributions of GIANT is the implementation of dynamic path feedback, which has been shown to enhance the learning efficiency. To visually demonstrate the efficacy of dynamic path feedback, we initially focus on GIANT’s reasoning ability on cluster-level, which directly impacts DWARF’s reasoning results. Here, we record the cluster state similarity (CSS) $D(s_{T}^{c},s_{\text{target}}^{c})$ under each epoch. As shown in Figure 5, our model outperforms CURL in scores and stability in most cases. This demonstrates that dynamic path feedback does improve GIANT’s learning efficiency. See Appendix B for a comparison of the remaining tasks.
Entity Reasoning Accuracy.  The preceding analysis has demonstrated that FULORA enables GIANT to provide high-quality guidance, and next, we will demonstrate that efficient guidance-exploration gives DWARF a superior performance in long-distance reasoning against other multi-hop baselines. Figure 6 plots the MRR scores on varying path lengths on NELL-995 and WN18RR. As can be seen, FULORA outperforms CURL, MINERVA and SQUIRE with different path lengths. In particular, FULORA is much more prominent in the long-distance reasoning accuracy of WN18RR (sparse KG). It explains our motivation: FULORA can alleviate the poor performance of multi-hop reasoning approaches in the absence of a short direct path (typically found in standard KG and sparse KG) by improving long-distance reasoning ability. Overall, FULORA has a more robust performance under the demands of long-distance reasoning, and also shows minimal performance degradation for long-path settings. It is mainly due to the dynamic path feedback and efficient guidance-exploration proposed by us, which enable more efficient information sharing between GIANT and DWARF, and ensure that DWARF maintains an excellent self-exploration ability.
Figure 6: The long-distance performance: FULORA significantly outpeforms CURL, MINERVA, SQUIRE on NELL-995 and WN18RR.
@@
We utilize $\delta$ and $\alpha$ to control the degree of efficient guidance-exploration and dynamic path feedback. To comprehensively explore the efficacy of efficient guidance-exploration and dynamic path feedback, we conduct link prediction on NELL-995 and WN18RR by varying the values of $\delta$ and $\alpha$ as $\{0.20,0.30,0.40,0.50\}$ and $\{0.05,0.10,0.15,0.20\}$, respectively. The averaged results are shown in Figure 7. In accordance with our previous analysis, it is evident that DWARF cannot rely excessively on GIANT for guidance. Furthermore, the level of path feedback that GIANT receives should not be unduly strong. In addition, in Appendix B, we have carefully discussed the influence of three main components of FULORA on knowledge graph reasoning efficiency.
@@ Conclusion
We present FULORA, an efficient guidance-exploration model built on dual-agent KG reasoning framework to enhance the agent’s long-distance reasoning ability on standard KG and sparse KG. The key insight behind our approach is balancing the self-exploration of DWARF and the guidance from GIANT. Specifically, on the one hand, we leverage the attention mechanism to make DWARF pay attention to the neighbouring entities that are close to the query. On the other hand, we propose that dynamic path feedback enables GIANT to have better learning efficiency, thus providing DWARF with high-quality guidance, making the DWARF to have a favourable global vision while having excellent local reasoning ability. Experiments on three real-world datasets show that FULORA outperforms the recent state-of-the-art multi-hop reasoning methods in most cases. Further analysis of these results reveals that FULORA’s Long-distance reasoning ability on standard KG and sparse KG has significant advantages compared to multi-hop reasoning methods.
Appendix A A. Experiment Setup
Data Statistics
We adopt three knowledge graph datasets with different scales: NELL-995, WN18RR and FB15K-237. We give a brief overview of these datasets.
NELL-995 (Xiong, Hoang, and Wang 2017). NELL995 is an open source machine learning dataset developed by the OpenAI research group that contains more than 950,000 pieces of entity relationship data collected from the network to help machine learning systems make inferences. NELL995 can be used to train machine learning models, such as natural language processing models, machine translation models, question answering systems, and semantic search systems.
WN18RR (Dettmers et al. 2018). WN18RR is a subset of WordNet that describes the association characteristics between English words. It preserves the symmetry, asymmetry, and composition relationships of the WordNet, and removes the inversion relationships. WN18RR contains some relational information about words. It consists of 40,945 entities and 11 relations.
FB15K-237 (Toutanova et al. 2015). FB15K-237 is a common knowledge graph dataset, which is a subset extracted from Freebase knowledge graph. It contains 14,505 entities and 237 relations, and the data is carefully processed to remove reversible relational data and trivial triples, ensuring that entities in the training set are not directly connected to the verification or test set, thus avoiding information leakage issues.
Baseline Methods
We compare the performance of FULORA with the following baselines, including embedding-based and multi-hop-based knowledge graph reasoning methods:
@@
    M-Walk (Shen et al. 2018) uses RNNs to convert the historical path into a vector for learning policy and Q-function. At the same time, because of the sparsity of reward, RNN with Monte Carlo tree search is used to generate paths.
@@
Implementation Details
We use the following software versions:
  • Ubuntu 24.04 LTS
  • Python 3.10
  • Pytorch 2.0.1
We conduct all experiments with a single NVIDIA GeForce 3080Ti GPU.
Our experiments of all of 10 baselines are conducted on their official implementation provided by their respective authors. To reproduce the results of our model in Table 2 and Table 3, we report the empirically optimal crucial hyperparameters as shown in Table 4. On all datasets, the quantities of path rollouts in training and testing are 20 and 100, separatively.
Attribute NELL-995 WN18RR FB15K-237
@@
Appendix B B. Additional Results and Analysis
Case Stuides

### v3

@@ Introduction
$\Delta(s_{t}^{c},s_{t+1}^{c})=\text{Sim}(s_{t}^{c},s_{\text{target}}^{c})-\text{Sim}(s_{t+1}^{c},s_{\text{target}}^{c})$
$s_{t+1}^{c}$ comes from the next state generated by the policy network. In contrast to the default reward, dynamic path feedback uses the reward function to score the GIANT’s path in a rollout rather than simply identifying whether it has reached the correct target cluster. Even if GIANT does not reach the correct target cluster in a rollout, it will evaluate the quality of the path, thereby accelerating the learning process. In Appendix D, we prove that GIANT learns optimal policy in dynamic path feedback is consistent with default rewards circumstance:
Theorem 1 (Consistency of optimal policy)
Given two MDPs that differ only in reward function, denoted as $M=(\mathcal{S},\mathcal{A},\delta,\mathcal{R})$ and $M^{\prime}=(\mathcal{S},\mathcal{A},\delta,\mathcal{R}_{D})$ respectively, where $\mathcal{R}=r_{c}(s_{t}^{c})$ is the default reward while $\mathcal{R}_{D}=r_{c}(s_{t}^{c})-\alpha\Delta(s_{t}^{c},s_{t+1}^{c})$ is the dynamic path feedback reward. Their optimal policies are consistent, that is,
$\pi^{*}_{M}(s_{t}^{c})=\pi^{*}_{M^{\prime}}(s_{t}^{c}).$ (10)
@@
Dataset #Ent #Rel #Fact #Que #Mean #Med
@@
Table 1: The statistics of some benchmark KG datasets. #Mean is the averaged outgoing degree of every entity that can indicate the sparsity level while #Med is the corresponding median.
@@
TransE (Bordes et al. 2013) 51.4 45.6 67.8 75.1 35.9 28.9 46.4 53.4 36.1 24.8 40.1 45.0
DistMult (Yang et al. 2015) 68.0 61.0 73.3 79.5 43.3 41.0 44.1 47.5 37.0 27.5 41.7 56.8
ComplEx (Trouillon et al. 2016) 68.4 61.2 76.1 82.1 41.5 38.2 43.3 48.0 39.4 30.3 43.4 57.2
LMKE (Wang et al. 2022) 74.6 71.7 84.7 89.5 53.6 45.1 66.0 79.4 41.2 31.8 46.2 56.9
NBFNet (Zhu et al. 2021) 70.9 68.0 76.8 80.4 48.1 44.9 49.8 58.3 40.5 30.8 45.8 55.2
RED-GNN (Zhang and Yao 2022) 71.2 67.8 - 86.2 46.9 42.5 - 53.0 39.8 29.9 - 54.4
MINERVA (Das et al. 2018) 67.5 58.8 74.6 81.3 44.8 41.3 45.6 51.3 27.1 19.2 30.7 42.6
AttnPath (Wang et al. 2019) 69.3 62.7 73.9 80.1 42.9 40.7 44.3 52.9 31.9 24.1 40.4 43.8
SQUIRE (Bai et al. 2022) 71.1 68.2 81.4 87.2 48.2 45.0 51.4 59.7 35.0 26.5 41.7 50.3
CURL (Zhang et al. 2022) 70.8 66.7 78.6 84.3 46.0 42.9 47.1 52.3 30.6 23.9 38.1 50.9
HMLS (Zheng et al. 2024) 71.8 69.0 80.9 88.9 48.5 43.9 52.9 60.4 37.4 28.0 42.1 52.5
FULORA (Ours) 72.5 69.4 79.7 89.2 49.1 45.6 50.7 59.2 36.4 27.1 41.9 51.3
@@
Table 2: Link prediction results with a path length of 3 on the NELL, WordNet, and Freebase datasets (embedding-based and GNN-based reasoning do not inherently consider path length, so we exclude triples reachable by shorter paths for fair comparison). All metrics are multiplied by 100. The best score of embedding-based reasoning models, GNN-based models are underlined while multi-hop reasoning models are in bold. Compared to other indicators, we specifically highlight the MRR to emphasize its significance.
As can be seen from the statistical indicators, these three datasets represent standard KG, sparse KG and Dense KG respectively. In our selection of the baseline, we are not only evaluating it against the state-of-the-art multi-hop reasoning methods, but also against with other embedding-based KG reasoning methods and GNN-based reasoning methods. FULORA and all baselines are implemented under the Pytorch framework and run on the NVIDIA 3080Ti GPU. All the results are the average of the results in five experiments. See Appendix A.1 for a brief introduction of each baseline.
Short-distance Reasoning Ability
For each triplet $(e_{s},r_{q},e_{o})$ in the test set, we convert it to a query $(e_{s},r_{q},?)$ and use embedding-based or multi-hop models with a beam search width of 50 to rank the tail entities. Following (Bordes et al. 2013), we evaluate using two metrics: (1) mean reciprocal rank (MRR) and (2) Hits@K, the proportion of correct tail entities ranked in the top K.
As shown in Table 2, we present the performance of FULORA and all baselines on NELL-995, WN18RR, and FB15K-237. On both the standard KG (NELL-995) and sparse KG (WN18RR), FULORA not only outperforms CURL, SQUIRE and HMLS currently regarded as the most powerful multi-hop reasoning models, but also significantly surpasses NBFNet and RED-GNN, both of which are recognized as advanced GNN-based reasoning algorithms. FULORA also achieves comparable performance to the best embedding-based model, LMKE. On FB15K-237, where 1-to-M relations dominate, multi-hop models often struggle with high-degree nodes, hindering correct entity retrieval. In contrast, FULORA’s attention mechanism focuses on the most relevant neighbors. Averaging across three datasets, FULORA improves MRR and Hits@1, 3, and 10 by 3.5%, 2.9%, 2.8%, and 4.1% over CURL. Notably, Table 2 presents results based on a path length of 3, which does not capture FULORA’s exceptional performance in long-distance reasoning. Subsequent experiments focus on the performance of FULORA’s components, GIANT and DWARF, in long-distance reasoning, with comparisons to other advanced models in Table 2. In fact prediction, FULORA outperforms other multi-hop baselines, as detailed in Appendix B.1.
NELL-995/WN18RR/FB15K-237 MRR @1 @10
FULORA + TransE (Bordes et al. 2013) 72.5/49.1/33.4 69.4/45.6/25.4 89.2/57.2/50.7
FULORA + LMKE (Wang et al. 2022) 76.8/54.4/38.6 73.0/46.3/30.2 93.7/72.4/51.2
Table 3: Effects of various pre-embedding methods on the performance of FULORA.
@@
Recall the motivation of FULORA, we tend to address the issue of multi-hop reasoning models being unable to infer correct answer due to the lack of short direct path by improving long-distance reasoning ability. This issue is significant on standard KG and sparse KG, so we conduct the following experiments on NELL-995 and WN18RR. For effective evaluation, we compare our model with MINERVA and CURL in NELL-995 and WN18RR, where we remove the most frequently-visited short paths found by the bi-directional search (Xiong, Hoang, and Wang 2017; Zhang et al. 2022) inside KGs.
@@
Cluster State Similarity.   One of the most significant contributions of GIANT is the implementation of dynamic path feedback, which has been shown to enhance the learning efficiency. To visually demonstrate the efficacy of dynamic path feedback, we initially focus on GIANT’s reasoning ability on cluster-level, which directly impacts DWARF’s reasoning results. Here, we record the cluster state similarity (CSS) $\text{Sim}(s_{T}^{c},s_{\text{target}}^{c})$ under each epoch. As shown in Figure 4, our model outperforms CURL in scores and stability in most cases. See Appendix B.3 for a comparison of the remaining tasks.
Entity Reasoning Accuracy.  we now demonstrate that FULORA’s efficient guidance-exploration enhances DWARF’s performance in long-distance reasoning compared to other well-performed baselines (Appendix B.4.2 provides a further analysis). Figure 5 shows the MRR for varying path lengths on NELL-995 and WN18RR. FULORA excels in long-distance reasoning accuracy on WN18RR (sparse KG). This highlights our motivation: FULORA mitigates the poor performance of current KG reasoning methods in the absence of short direct paths by enhancing long-distance reasoning. Overall, FULORA demonstrates more robust long-distance reasoning performance with minimal degradation in long-path settings. This is due to the dynamic path feedback and efficient guidance-exploration, which enable better information sharing between GIANT and DWARF, ensuring DWARF’s excellent self-exploration ability.
Figure 5: The long-distance performance: FULORA significantly outpeforms CURL, SQUIRE, HMLS, LMKE, RED-GNN, NBFNet on NELL-995 (standrad KG) and WN18RR (Sparse KG).
@@
Sensitivity Analysis.  We utilize $\delta$ and $\alpha$ to control the degree of efficient guidance-exploration and dynamic path feedback. To comprehensively explore the efficacy of the two mechanisms, we conduct link prediction on NELL-995 and WN18RR by varying the values of $\delta$ and $\alpha$ as $\{0.20,0.30,0.40,0.50\}$ and $\{0.05,0.10,0.15,0.20\}$, respectively. The averaged results are shown in Figure 6. In accordance with our previous analysis, it is evident that DWARF cannot rely excessively on GIANT for guidance. Furthermore, the level of path feedback that GIANT receives should not be unduly strong. In addition, in Appendix B.4, we have carefully discussed the influence of three main components of FULORA on KG reasoning efficiency.
@@
Pre-embedding Methods.  In our approach, we use TransE (Bordes et al. 2013) for pre-training the embeddings. One straightforward idea is to improve FULORA’s performance by using a more advanced KG embedding model such as LMKE (Wang et al. 2022). Table 3 presents the performance of TransE and LMKE as pre-embedding methods. The more advanced LMKE indeed improves FULORA’s performance. However, we use TransE in our experiments to emphasize that the performance improvement is due to FULORA’s efficient guidance-exploration mechanism, not the strength of the KG embedding method.
@@ Conclusion
We present FULORA, an efficient guidance-exploration model built on dual-agent KG reasoning framework to enhance the agent’s long-distance reasoning ability on standard KG and sparse KG. The key insight behind our approach is balancing the self-exploration of DWARF and the guidance from GIANT. Specifically, on the one hand, we leverage the attention mechanism to make DWARF pay attention to the neighbouring entities that are close to the query. On the other hand, we propose that dynamic path feedback enables GIANT to have better learning efficiency, thus providing DWARF with high-quality guidance, making the DWARF to have a favourable global vision while having excellent local reasoning ability. Experiments on three real-world datasets demonstrate that FULORA outperforms state-of-the-art multi-hop reasoning methods. Further analysis reveals that FULORA’s long-distance reasoning ability on standard and sparse KGs significantly outperforms current KG reasoning methods.
Appendix A A  Experiment Setup
A.1  Baseline Methods
We compare the performance of FULORA with the following baselines [^2] Code to reproduce the main experiments is available at: https://github.com/KotoHanon/LOCOCO., including embedding-based, GNN-based and multi-hop based KG reasoning methods:
@@ A.1  Baseline Methods
    NBFNet (Zhu et al. 2021) parameterizes the generalized Bellman-Ford algorithm with 3 neural components, namely INDICATOR, MESSAGE and AGGREGATE functions, which corresponds to the boundary condition, multiplication operator, and summationoperator respectively
@@
    RED-GNN (Zhang and Yao 2022) makes use of dynamic programming to recursively encodes multiple r-digraphs with shared edges, and utilizes query-dependent attention mechanism to select the strongly correlated edges.
  9. 9.
@@
  12. 12.
    HMLS (Zheng et al. 2024) improves the generalizability and effectiveness of multi-hop reasoning in few-shot scenarios by exploiting hard relations and hierarchical relation structures.
Hyperparameter NELL-995 WN18RR FB15K-237
@@
Task TransE TransR PRA DeepPath MINERVA M-Walk CURL FULORA
PersonBornInLocation 62.7 67.3 54.7 75.1 80.0 84.7 82.7 84.4
OrgHeadquarteredInCity 62.0 65.7 81.1 79.0 94.0 94.3 94.8 95.1
AthletePlaysForTeam 62.7 67.3 54.7 75.0 80.3 84.7 82.9 86.0
AthletePlaysInLeague 77.3 91.2 84.1 96.0 94.2 96.1 97.1 97.3
AthletePlaysSport 87.6 96.3 47.4 95.7 98.0 98.3 98.4 98.5
TeamPlaysSport 76.1 81.4 79.1 73.8 88.0 88.4 88.7 90.2
WorksFor 67.7 69.2 68.1 71.1 81.0 83.2 82.1 84.3
A.2  Data Statistics
We adopt three KG datasets with different scales: NELL-995 (standard KG), WN18RR (sparse KG) and FB15K-237 (dense KG). We give a brief overview of these datasets.
NELL-995 (Xiong, Hoang, and Wang 2017) is an open source machine learning dataset developed by the OpenAI research group that contains more than 950,000 pieces of entity relationship data collected from the network to help machine learning systems make inferences. NELL995 can be used to train machine learning models, such as natural language processing models, machine translation models, question answering systems, and semantic search systems.
WN18RR (Dettmers et al. 2018) is a subset of WordNet that describes the association characteristics between English words. It preserves the symmetry, asymmetry, and composition relationships of the WordNet, and removes the inversion relationships. WN18RR contains some relational information about words. It consists of 14,541 entities and 237 relationships.
FB15K-237 (Toutanova et al. 2015) is a common knowledge graph dataset, which is a subset extracted from Freebase knowledge graph. It contains 14,505 entities and 237 relationships, and the data is carefully processed to remove reversible relational data and trivial triples, ensuring that entities in the training set are not directly connected to the verification or test set, thus avoiding information leakage issues.
Appendix B B  Additional Results and Analysis
B.1  Fact Prediction Results
As opposed to link prediction, fact prediction task is concerned with verifying the veracity of an unknown fact, the true test triplets are ranked with some generated false triplets. Since we share a similar query-answering mechanism as CURL (Zhang et al. 2022), FULORA is capable of identifying the most appropriate entity for a given query and eliminates the need to evaluate negative samples of any particular relation. In the experiments, the dual agents try to infer and walk through the cluster-level and entity-level respectively to reach the correct target under removing all links of groundtruth relations in the original KG. Here, we report Mean Average Precision (MAP) scores for various relation tasks of NELL-995. We reuse the results of TransE (Bordes et al. 2013), TransR (Lin et al. 2015), PRA (Lao and Cohen 2010), DeepPath (Xiong, Hoang, and Wang 2017), MINERVA (Das et al. 2018), M-Walk (Shen et al. 2018) on seven tasks already reported in (Zhang et al. 2022). As demonstrated in Table 5, FULORA produces a satisfying result in most tasks, contributing an average gain of 9.1% relative to the multi-hop based reasoning approaches (PRA, DeepPath, MINERVA, M-Walk and CURL) and 16.1% gain compared to the embedding-based approaches (TransE and TransR).
B.2  Case Stuides